# IP-CRR: Information Pursuit for Interpretable Classification of Chest Radiology Reports

### v1

@@ Abstract
The development of AI-based methods for analyzing radiology reports could lead to significant advances in medical diagnosis—from improving diagnostic accuracy to enhancing efficiency and reducing workload. However, the lack of interpretability in these methods has hindered their adoption in clinical settings. In this paper, we propose an interpretable-by-design framework for classifying radiology reports. The key idea is to extract a set of most informative queries from a large set of reports and use these queries and their corresponding answers to predict a diagnosis. Thus, the explanation for a prediction is, by construction, the set of selected queries and answers. We use the Information Pursuit framework to select informative queries, the Flan-T5 model to determine if facts are present in the report, and a classifier to predict the disease. Experiments on the MIMIC-CXR dataset demonstrate the effectiveness of the proposed method, highlighting its potential to enhance trust and usability in medical AI. Code is available at: https://github.com/Glourier/IP-CRR.git.
@@
Interpretable machine learning Text classification Radiology reports.
@@ 1 Introduction
When a doctor makes a decision that can potentially have life-changing consequences for a patient, the transparency of the doctor’s decision-making is invaluable. Therefore, as machine learning methods become increasingly used in clinical practice, there is a pressing need to develop safe and trustworthy models. For a given prediction task, interpretability in machine learning serves as a crucial aspect of model transparency that answers the “why’s” and “how’s” of a model’s prediction, fostering trust between the user and the machine.
@@
Traditionally, interpretability is carried out by first training a black-box, opaque model that optimizes performance, and then producing an explanation in the form of heatmaps [30, 6], selection of features [28, 21], or more generally, attributions [33, 31]. These post-hoc explanation methods have found overwhelming success in many areas such as computer vision, finance, and even healthcare. However, these methods lack faithfulness [29]. That is, the explanation could potentially be an artifact of the explanation method rather than evidence of the model’s underlying reasoning. To address this, interpretable-by-design models were proposed to make predictions solely based on a set of interpretable factors [16]. Interpretable factors take the form of a sparse set of concepts [18] or a set of questions and answers [7] about the given test sample. For example, disease diagnosis for a patient can be conducted by a list of symptoms, which indicates the explanations for the final prediction. While interpretable-by-design methods address the issue of faithfulness, they often require making careful and meticulous designs about the model architecture, optimization, and the set of interpretable concepts or questions. This complexity poses a significant challenge, particularly in the medical domain, where the need for both accuracy and interpretability is critical for clinical decision-making.
Figure 1: Example of the interpretable-by-design method.
@@
Take the task of Chest Radiology Report (CRR) classification as an example (Figure 1). A given report typically contains a few key insights about the patient, and the task is to predict the disease that the patient presents. For the prediction to be interpretable, it must be grounded on a set of queries about the patient or the report, such as “Is the heart contour enlarged?” or “Does the lung volume decrease?”. To be able to make a prediction, we need not only the queries, but also the answers. Importantly, there might potentially exist queries that do not have readily available answers in the report. This setting drastically differs from the existing settings that interpretable-by-design methods are applied to, where all answers are typically available and easily obtained using out-of-the-shelf models [8].
@@
In this work, we propose an interpretable-by-design framework for CRR classification, which extends the Information Pursuit (IP) framework for explainable classification of natural images [9, 7]. Our proposed IP-CRR method involves: 1) curating a set of interpretable queries by mining them from existing radiology reports; 2) applying natural language inference to determine if a fact is present in the report to be classified; and 3) learning a classifier that makes predictions by sequentially selecting a list of queries in order of information gain. Implementing our approach requires two key innovations relative to [9, 7]: 1) modifying the IP framework to accommodate large language models (LLMs) for answering queries, and 2) handling “Unknown” answers, as illustrated in Figure 1. Empirically, our method adapts to queries with readily available answers in the report by leveraging natural language inference. Moreover, our method achieves higher F1 score against black-box language baselines such as Flan-T5-large as well as interpretable-by-design methods such as Concept Bottleneck Models. Qualitatively, the explanations provided by IP-CRR offer a potential venue for trustworthy and transparent predictions in clinical decision-making.
@@ 2 Related Work
Radiology Report Classification. Automatic disease diagnosis from radiology reports is a critical task that enables large-scale use of machine learning models in clinical settings. Natural language processing methods are widely used to extract diagnostic labels from free-text radiology reports. There are two main approaches for this task: rule-based labeling and machine learning-based labeling. Rule-based methods [5, 25, 4, 14] rely on predefined rules or regular expressions to extract diseases from the report. However, these methods struggle with the complexity and variability of clinical language. Machine learning-based methods [32, 34, 26, 13] learn the patterns of text automatically without relying on handcrafted rules. For example, CheXbert [32] is a BERT-based medical domain specialized model pretrained on rule-based annotations, and finetuned on a small expert annotated dataset. However, despite their performance, they lack interpretability, making it hard for radiologists to trust their predictions in clinical use.
Interpretable-by-design Methods. Interpretable-by-design methods aim to produce explanations that are faithful to the model. State-of-the-art methods ensure this by using an intermediate layer of interpretable factors (e.g., concepts or questions) based on which a prediction is made. For example, Concept Bottleneck Models (CBMs) first map a given input to a list of concept scores, each corresponding to the presence of a concept in the input, and then linearly map concept scores to a predicted label. CBMs have found many successful applications in domains such as image classification and clinical prediction [23, 10, 35, 17, 24]. However, CBMs are sensitive to having concepts of very high quality and their use of a linear classifier limits their accuracy. In this work, we focus on another framework known as Information Pursuit (IP) [9, 7, 12], which provides interpretability via sequentially selecting queries in order of information gain, and then making predictions based on the sequence of query-answer pairs. Therefore, IP aligns more closely with the interactive question-answering process between a doctor and a patient. The challenge, however, involves constructing a set of meaningful queries and obtaining their answers. Currently, it has only been explored for the image classification domain, where [8] leverages large language models [1] and vision-language models [19]. In this work, we aim to leverage natural language inference for text and medical domain. We discuss the framework in detail in Section 3.3.
@@ 3 Method
In this section, we describe IP-CRR, the proposed interpretable-by-design model for classifying radiology reports. Our pipeline consists of three main components: 1) generate a set of queries from a large amount of reports (Sec. 3.2), 2) design a mechanism to answer them automatically (Sec. 3.2), and 3) integrate the Information Pursuit framework (Sec. 3.3) to conduct interpretable classification.
@@ 3.1 Problem Formulation
We consider the problem of classifying a radiology report based on a sequence of interpretable queries. Let $X\in\mathcal{X}$ be a random variable representing a radiology report, where $\mathcal{X}$ is the space of all possible reports, and $Y\in\mathcal{Y}=\{0,1\}$ denote the corresponding binary label. Our goal is to determine $Y$ given $X$ by sequentially querying $X$ with a set $Q$ of predefined, interpretable queries.
@@
Each query $q\in Q$ is a function $q:\mathcal{X}\to\mathcal{A}$, where $\mathcal{A}=\{-1,0,1\}$ represents the possible answers: negative, unknown, and positive. For example, given a radiology report $x$ and a query $q=$Is there a decreased opacity on the right?”, $q(x)=1$ implies there is a decreased opacity on the right, $q(x)=-1$ implies there is not, and $q(x)=0$ implies the query cannot be answered from the given report. We assume the query set $Q$ is interpretable, task-specific, and sufficient [^1] That is, $P(Y\mid x)=P(Y\mid\{x^{\prime}\in\mathcal{X}:q(x)=q(x^{\prime}),q\in Q\})$. See [9]. .
@@
The classification process involves selecting the most informative queries in an iterative manner until sufficient information is obtained to predict $Y$ with high confidence.
3.2 Constructing and Answering Queries from Existing Database
@@
Our final goal is to classify whether a certain disease is present or not in a given CRR $x$. As a prerequisite to applying the IP framework, we first need to define a set of task-relevant and interpretable queries that inform us about the report. In this work, we choose a query that corresponds to a question on whether a particular fact is present in the given report. For example, the fact “The heart shape is normal” can be thought of as a query by rephrasing it as “Did the report mention that the heart shape is normal?” While there are many potential choices for a query set, we argue that curating a sufficiently large set of facts as queries for CRR classification would allow us to build interpretable models with good predictive performance. In what follows, we describe how to construct this query set and how to answer queries for a given radiology report.
@@
Queries $q$. The procedure for obtaining a large set of representative, domain-specific, and interpretable queries is illustrated in Figure 2 Part I. We construct the query set $Q$ by mining facts from a large-scale radiology report dataset using a recently developed fact extraction model [22]. Specifically, we analysis over 220k radiology reports from MIMIC-CXR dataset [15], parse them into over 660k sentences using Natural Language Toolkit [2], and then extract over 590k facts from these sentences. To extract facts from sentences, we first prompt the GPT-3.5 and GPT-4 models to generate facts from each sentence on a subset of sentences, then use these input-output pairs as training data to finetune a sequence-to-sequence model named T5-small [27] for fact extraction. After fact extraction, we encode each fact into a 128-dimension latent space using the fact encoder CXRFE [22]. We apply $k$-means clustering to all 590k extracted facts in the fact embedding space, then select the nearest fact to each cluster center, and conduct deduplication if any two selected facts have a cosine similarity greater than 0.97. After this process, we obtain 520 representative facts, where each of them is a short fact describing the clinical findings, e.g., “moderate size hiatal hernia” or “increased bibasilar opacities”.
@@
Query Answers $q(x)$. Ideally, we hope to directly obtain grounded query answers from the report via annotations from radiologists. However, manual annotation is infeasible due to the large number of reports and queries. Moreover, in contrast to image classification, it often occurs that a query cannot be answered based on the provided radiology report. For example, if the description of the heart is not mentioned in the report, we cannot answer the query “Is the heart size normal?” from the report, although the heart is shown in the image. To obtain the query answers in a scalable and accurate manner, we propose to leverage natural language inference (NLI). Specifically, we pass both the query and the CRR as inputs to an LLM named Flan-T5-large [11] and ask the LLM to predict the query answer as either “yes”, “no” or “unknown”. An example of the LLM prompt is illustrated in Figure 2 Part II. The prompt consists of three parts: instruction, premise, and hypothesis. The instruction serves as guidance to the model and is fixed for every sample, the premise is the radiology report, and the hypothesis is the fact to be verified whether it is true for the given report or not. For every report and query, we set the predicted output of the model as the query answer $q(x)\in\{-1,0,1\}$.
@@ 3.3 Classification Based on Variational Information Pursuit
IP. We classify a report using Information Pursuit (IP), an interpretable-by-design framework that sequentially selects queries about the data in order of information gain. Given a sample $x^{\text{obs}}\in\mathcal{X}$, the IP algorithm selects the queries as follows:
@@
Here $q_{k+1}\in Q$ refers to the new query selected at step $k+1$ based on the history $q_{1:k}(x^{\text{obs}})$. The algorithm terminates at iteration $L$ when there are no more informative queries, i.e., $\forall q\in Q,I(q(X);Y|q_{1:L}(x^{\text{obs}}))=0$. Furthermore at each iteration $k$, one can also compute the posterior distribution (i.e. the prediction) $P(Y\mid q_{1:k}(x^{\text{obs}}))$. To carry out IP, [9] proposed to learn a VAE to model the joint distribution of all query answers and labels $P(Q(X),Y)$ directly. However, learning a good generative model has shown to be challenging when the size of $Q$ is large, so a variational approach is proposed as an alternative.
V-IP. Rather than modeling the joint distribution of queries and labels to directly compute the most informative queries, Variational Information Pursuit (V-IP) [7, 8, 12] finds a querier $g$ that selects the most informative query $q$ given any random history $S$ of query-answer pairs and a predictor $f$ that models the posterior distribution of $Y$ given $q$ and $S$, by optimizing the following objective:
$\min_{f,g}$ $\quad\mathbb{E}_{X,S}\left[D_{KL}(P(Y\mid X)\ \|\ \hat{P}(Y\mid q(X),S))\right]$ (2)
where $\quad q:=g(S)\in Q$ (3)
$\quad\hat{P}(Y\mid q(X),S):=f(\{q,q(X)\}\cup S).$ (4)
@@
To implement the whole framework of IP-CRR, we first generate a query set which consists of queries from large amounts of radiology reports, then answer each query for each report with an LLM. Finally, classification is conducted by generating a sequence of interpretable query-answers.
@@ 4 Experiments
We refer the entire pipeline, from generating queries from the radiology report database, answering queries via natural language inference, and applying V-IP to our task, as IP-CRR. Our proposed framework is modified to handle query answers $q(x)$ obtained from NLI, and our framework naturally handles the issue with unknown query answers. In this section, we will demonstrate its effectiveness in performing a range of tasks in radiology report classification.

### v2

@@ Abstract
The development of AI-based methods to analyze radiology reports could lead to significant advances in medical diagnosis, from improving diagnostic accuracy to enhancing efficiency and reducing workload. However, the lack of interpretability of AI-based methods could hinder their adoption in clinical settings. In this paper, we propose an interpretable-by-design framework for classifying chest radiology reports. First, we extract a set of representative facts from a large set of reports. Then, given a new report, we query whether a small subset of the representative facts is entailed by the report, and predict a diagnosis based on the selected subset of query-answer pairs. The explanation for a prediction is, by construction, the set of selected queries and answers. We use the Information Pursuit framework to select the most informative queries, a natural language inference model to determine if a fact is entailed by the report, and a classifier to predict the disease. Experiments on the MIMIC-CXR dataset demonstrate the effectiveness of the proposed method, highlighting its potential to enhance trust and usability in medical AI. Code is available at: https://github.com/Glourier/MICCAI2025-IP-CRR.
@@
Interpretable ML Text classification Radiology reports.
@@ 1 Introduction
When a doctor makes a decision that can potentially have life-changing consequences for a patient, the transparency of the decision-making is invaluable. As machine learning (ML) methods become more integrated into healthcare, there is a pressing need for safe and trustworthy models. For a prediction task, interpretability serves as a crucial aspect of model transparency that answers “why” and “how” a model made a prediction, fostering user trust.
@@
Traditionally, interpretability is carried out by first training a black-box model that optimizes performance, and then producing an explanation in the form of heatmaps [30, 6], selection of features [28, 21], or more generally, attributions [33, 31]. These post-hoc explanation methods are widely used in many areas such as computer vision, finance, and healthcare. However, these methods lack faithfulness [29], because the explanation could potentially be an artifact of the explanation method rather than evidence of the model’s underlying reasoning. To address this, interpretable-by-design models were proposed to make predictions solely based on a set of interpretable factors [16], which usually take the form of a set of concepts [18] or queries [9] about the input. For example, disease diagnosis for a patient can rely on a list of symptoms, which naturally explain the final prediction. While these methods improve faithfulness, they often require making careful and meticulous designs about the model architecture, optimization, and the set of interpretable concepts or questions. This complexity poses a significant challenge, particularly in the medical domain, where the need for both accuracy and interpretability are critical.
Figure 1: Illustration of the interpretable-by-design method.
@@
Take the task of Chest Radiology Report (CRR) classification as an example (Figure 1). A report typically contains a few key insights about the patient, and the task is to predict the associated diseases. For predictions to be interpretable, they must be grounded on a set of queries about the report or patient, such as “Is the heart contour enlarged?” or “Does the lung volume decrease?”. This requires both the queries and their corresponding answers. Importantly, some queries may lack readily available answers in the report. This setting drastically differs from most settings in interpretable-by-design methods, where all answers are typically available and easily obtained using out-of-the-shelf models [8].
@@
In this work, we propose an interpretable-by-design framework for CRR classification, which extends the Information Pursuit (IP) framework from image domain [9, 7] to text. Our proposed IP-CRR method involves: 1) curating a set of interpretable queries by mining them from existing CRRs; 2) applying natural language inference (NLI) to determine if a fact is entailed by the report; and 3) learning a classifier that makes predictions by sequentially selecting a list of queries in order of information gain, answering those queries, and predicting the class from the sequence of query-answer pairs. Implementing our approach requires two key innovations relative to [9, 7]: 1) modifying the IP framework to accommodate large language models (LLMs) for answering queries, and 2) handling “Unknown” answers, as illustrated in Figure 1. Moreover, our method achieves higher F1 score against black-box baselines such as Flan-T5-large and interpretable-by-design methods such as Concept Bottleneck Models. Qualitatively, the explanations provided by IP-CRR offer a potential venue for trustworthy and transparent predictions in clinical decision-making.
@@ 2 Related Work
Radiology Report Classification. Automatic disease diagnosis from radiology reports is a critical task that enables large-scale use of ML models in clinical settings. Natural language processing methods are widely used to extract diagnostic labels from free-text radiology reports. There are two main approaches for this task: rule-based labeling and ML-based labeling. Rule-based methods [5, 25, 4, 14] rely on predefined rules or regular expressions to extract diseases from the report. However, these methods struggle with the complexity and variability of clinical language. ML-based methods [34, 32, 26, 13] learn the patterns of text automatically without relying on handcrafted rules. For example, CheXbert [32] is a BERT-based medical domain specialized model pretrained on rule-based annotations, and finetuned on a small expert annotated dataset. However, despite their performance, they lack interpretability, making it hard for radiologists to trust their predictions in clinical use.
Interpretable-by-design Methods. Interpretable-by-design methods aim to produce explanations that are faithful to the model. State-of-the-art methods ensure this by using an intermediate layer of interpretable factors (e.g., concepts or queries) based on which a prediction is made. For example, Concept Bottleneck Models (CBMs) first map a given input to a list of concept scores, each corresponding to the presence of a concept in the input, and then linearly map concept scores to a predicted label. CBMs have found many successful applications in domains such as image classification and clinical prediction [23, 10, 35, 17, 24]. However, CBMs are sensitive to having concepts of very high quality and their use of a linear classifier limits their accuracy. In this work, we focus on another framework known as Information Pursuit (IP) [9, 7, 12], which provides interpretability via sequentially selecting queries in order of information gain, and then making predictions based on the sequence of query-answer pairs. Therefore, IP aligns more closely with the interactive question-answering process between a doctor and a patient. The challenge, however, involves constructing a set of meaningful queries and obtaining their answers. Currently, it has only been explored for image classification, where [8] leverages LLMs [1] and vision-language models [19]. In this work, we aim to leverage natural language inference for text and medical domain. We discuss the framework in detail in Section 3.3.
@@ 3 Method
In this section, we describe IP-CRR, the proposed interpretable-by-design model for classifying CRRs. Our pipeline consists of three main components: 1) generate a set of queries from a large number of reports (Sec. 3.2), 2) design a mechanism to answer queries automatically (Sec. 3.2), and 3) integrate the Information Pursuit framework (Sec. 3.3) to conduct interpretable classification.
@@ 3.1 Problem Formulation
We consider the problem of classifying a CRR based on a sequence of interpretable queries. Let $X\in\mathcal{X}$ be a random variable representing a CRR, where $\mathcal{X}$ is the space of all possible CRRs, and $Y\in\mathcal{Y}=\{0,1\}$ denotes the corresponding binary label. Our goal is to determine $Y$ given $X$ by sequentially querying $X$ with a set $Q$ of predefined, interpretable queries.
@@
Each query $q\in Q$ is a function $q:\mathcal{X}\to\mathcal{A}$, where $\mathcal{A}=\{-1,0,1\}$ represents the possible answers: negative, unknown, and positive. For example, given a radiology report $x$ and a query q=q=Is there a decreased opacity on the right?”, $q(x)=1$ implies there is a decreased opacity on the right, q(x)=1q(x)=-1 implies there is not, and $q(x)=0$ implies the query cannot be answered from the given report. We assume the query set $Q$ is interpretable, task-specific, and sufficient [^1] That is, $P(Y\mid x)=P(Y\mid\{x^{\prime}\in\mathcal{X}:q(x)=q(x^{\prime}),q\in Q\})$. See [9]. .
@@
The classification process involves selecting the most informative queries iteratively until sufficient information is obtained to predict $Y$ with high confidence.
3.2 Constructing and Answering Queries from Existing Databases
@@
Our final goal is to classify whether a certain disease is present or not in a given CRR. To apply the IP framework, we first need to define a set of interpretable and task-relevant queries about the report. Each query corresponds to a question about whether a specific fact is “true”, “false”, or “unknown” given the report. For example, the fact “normal heart shape” can be thought of as the query “Is the heart shape normal?” While there are many choices for the query set, we argue that curating a sufficiently large set of facts as queries for CRR classification allows us to build interpretable models with good predictive performance. Next, we describe how to construct this query set and answer queries for a CRR.
@@
Queries $q$. The procedure for obtaining a large set of representative, domain-specific, and interpretable queries is illustrated in Figure 2 Part I. We construct the query set $Q$ by mining facts from a large-scale CRR dataset using a recently developed fact extraction model [22]. Specifically, we analyze over 220k CRRs from the MIMIC-CXR dataset [15] and parse them into more than 660k sentences using the Natural Language Toolkit [2]. Next, we extract over 590k facts from these 660k sentences by first prompting the GPT-3.5 and GPT-4 models to generate facts for a subset of the sentences and then using these sentence-fact pairs as training data to finetune the T5-small [27] sequence-to-sequence model for fact extraction. After fact extraction, we encode each fact into a 128-dimension latent space using the fact encoder CXRFE [22]. We apply $k$-means clustering to all 590k extracted facts in the fact embedding space, then select the nearest fact to each cluster center, and conduct deduplication if any two selected facts have a cosine similarity greater than 0.97. After this process, we obtain 520 representative facts, where each of them is a short fact describing the clinical findings, e.g., “moderate size hiatal hernia” or “increased bibasilar opacities”.
@@
Query Answers $q(x)$. Ideally, query answers should be obtained from the report through annotations made by radiologists. However, manual annotation is infeasible due to the large number of reports and queries. Moreover, in contrast to image classification, it often occurs that a query cannot be answered based on the provided radiology report. For example, if the description of the heart is not mentioned in the report, we cannot answer the query “Is the heart size normal?” from the report, although the heart is shown in the image. To obtain the query answers in a scalable and accurate manner, we propose to leverage natural language inference (NLI). Specifically, we pass both the query and the CRR as inputs to the Flan-T5-large [11] LLM and ask it to predict the query answer as either “yes”, “no” or “unknown”. An example of the prompt is illustrated in Figure 2 Part II. The prompt consists of three parts: instruction, premise, and hypothesis. The instruction serves as guidance to the model and is fixed for every sample, the premise is the radiology report, and the hypothesis is the fact to be verified as true for the given report. For every report and query, the original model output is 0, 1, or 2 for negative, unknown, or positive. Then we normalize them to form the query answer $q(x)\in\{-1,0,1\}$.
@@ 3.3 Classification Based on Variational Information Pursuit
IP. We aim to classify a report using Information Pursuit (IP), an interpretable-by-design framework that sequentially selects queries about the data in order of information gain. Let $I(q(X);Y)$ denote the mutual information between $q(X)$ and $Y$. Given a sample CRR $x^{\text{obs}}\in\mathcal{X}$, IP selects the queries as follows:
@@
Here, $q_{k+1}\in Q$ refers to the new query selected at step $k+1$ based on the history of query-answer pairs $q_{1:k}(x^{\text{obs}})$. The algorithm terminates at iteration $L$ when there are no more informative queries, i.e., $\forall q\in Q,I(q(X);Y\mid q_{1:L}(x^{\text{obs}}))=0$. Furthermore at each iteration $k$, one can also compute the posterior distribution (i.e. the prediction) $P(Y\mid q_{1:k}(x^{\text{obs}}))$. To carry out IP, [9] proposed to learn a VAE to model the joint distribution of all query answers and labels $P(Q(X),Y)$. However, learning a good generative model is challenging when the size of $Q$ is large, so a variational approach is proposed as an alternative.
V-IP. Rather than modeling the joint distribution of queries and labels to directly compute the most informative queries, Variational Information Pursuit (V-IP) [7, 8, 12] finds a querier $g$ that selects the most informative query $q$ given any random history $S$ of query-answer pairs, and a predictor $f$ that models the posterior distribution of $Y$ given $q$ and $S$, by optimizing the following objective:
$\begin{split}\min_{f,g}&\quad\mathbb{E}_{X,S}\left[D_{KL}(P(Y\mid X)\ \|\ \hat{P}(Y\mid q(X),S))\right]\\ \text{where}&\quad q:=g(S)\in Q\\ &\quad\hat{P}(Y\mid q(X),S):=f(\{q,q(X)\}\cup S).\end{split}$ (V-IP)
@@
To implement IP-CRR, we first obtain a query set from large amounts of CRRs as described in Sec. 3.2, then answer each query for each report with an LLM. Finally, we apply V-IP to train a querier and a classifier to generate a sequence of interpretable query-answer pairs for the final classification.
@@ 4 Experiments
In this section, we will demonstrate the effectiveness of IP-CRR at performing CRR classification.